Suppose you travel north for 65 kilometers then travel east 75 kilometers. How far are you from your starting point? show steps taken.
The volume of a cube is given by V = s3, where s is the length of a side. Find the length of a side of a cube if the Volume is 1000 cm3.show steps taken.

Answer by Nate(3500)

Suppose you travel north for 65 kilometers then travel east 75 kilometers. How far are you from your starting point? show steps taken.
Draw a vertical line and label it 65.
Draw another vertical line making a 90-degree angle and label it 75.
Connect the two lines with a vertical line and label it x.
... You should get a right triangle.
x^2 = 65^2 + 75^2
x^2 = 4225 + 5625
x^2 = 9850
About 99.25km
~
The volume of a cube is given by V = s3, where s is the length of a side. Find the length of a side of a cube if the Volume is 1000 cm3.show steps taken.
V = s^3
1000 = s^3
1000^(1/3) = s
(10*10*10)^(1/3) = s
10cm = s